OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price

OpenStudy (anonymous):

Trying to figure out the price in september.

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Ratio R = September price : October price = 100% : 120% = x : $1.50 so we have R=100/120=x/$1.50 To find the September price Solve for x in \(\Large \frac{100}{120}=\frac{x}{150}\)